discount sound has 260 retail


1
Expert's answer
2021-07-15T10:22:07-0400

We know that the margin of the error "=z_{\alpha/2}\frac{(\sigma)}{\sqrt{n}}"


"=\frac{1.96\times 4500}{\sqrt{36}}"


"=1.96\times 750"


"=1470"

Interval estimate of "\mu =\$31000\pm 1470"

Hence, "\$ 29630" to the "\$32570"